4-2y=10-6y what does 'y' equal?
Answer by queenofit(120)

First you must get all your "team members" to the correct side of the = sign.
(put the variable "team" on the other side of the = sign) remember when a number or variable changes sides, you have to change its value from positive to negitive and vise versa. It doesn't matter which way you go, but I prefer to make the variable positive.
4-2y=10-6y
would become
6y-2y=10-4 (did you see how I changed the +/- signs when I moved the equation)
4y=6 You must get rid of the 4 because it is not on the Variable "team" You
divide each side by 4 to get rid of the number in front of y.
y= 6/4 depending on what type of answer your instructor wants, change the 6/4.
For example, my instructor always said this was far enough in college, but in high school they wanted us to reduce 6/4 = 3/2 or 1 1/2 . Some instructors want the decimal version of the answer. This is completely up to your instructor for his/her preference. If you need a decimal answer, divide 6 by 4 = 1.5
Y= 6/4 or y = 3/2 or y= 1 & 1/2 or y= 1.5 One of these will be in the form your instructor likes. You are the only one who knows how they like you to show your answer...